jagr2713 (jagr2713):

f-1(3) when f(x) = 2x plus 3, all over 5

jagr2713 (jagr2713):

@iambatman @jim_thompson5910 @radar @Zarkon @jdoe0001

jagr2713 (jagr2713):

And please explain your steps

OpenStudy (anonymous):

Alright, we want the inverse?

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x) Follow these steps, and see what you get, and once you have the inverse plug in the 3.

jagr2713 (jagr2713):

@iambatman i dont get it can you do it please?

OpenStudy (anonymous):

You didn't even try it.

jagr2713 (jagr2713):

i did that before i dont get it

OpenStudy (anonymous):

I'm not going to do this for you, I will help guide you but you need to do the heavy lifting.

OpenStudy (anonymous):

Well can you show me your process?

OpenStudy (anonymous):

I can then pin point where you went wrong.

jagr2713 (jagr2713):

like i replaced X with y

OpenStudy (anonymous):

\[f(x) = \frac{ 2x+3 }{ 5 } \implies y = \frac{ 2x+3 }{ 5 }\]

jagr2713 (jagr2713):

yea that and i was stuck after that

OpenStudy (anonymous):

Follow the steps I gave you, switch the x's and y's.

jagr2713 (jagr2713):

so x=2y+3/5

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Now solve for y

jagr2713 (jagr2713):

then we multiply right?

OpenStudy (anonymous):

Multiply what?

jagr2713 (jagr2713):

5 and 3?

OpenStudy (anonymous):

Yes you multiply 5 but not 3, write it out, it will make much more sense.

jagr2713 (jagr2713):

so u muliply 5 with x and get 5x=2y+3

jagr2713 (jagr2713):

@iambatman

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What you do to one side you must do to the other, remember this, so do same stuff on your right side on the left as well. That's algebra.

jagr2713 (jagr2713):

then we reverse the equality and get 2y+3=5x?

OpenStudy (anonymous):

It doesn't matter, but that's fine.

OpenStudy (anonymous):

Solve for y, what's your next step/

jagr2713 (jagr2713):

then isolate and get 2y=5x-3 then solve for y >?

jagr2713 (jagr2713):

so its y=5x/2 - 3/2

OpenStudy (anonymous):

Yes that's good

jagr2713 (jagr2713):

So i am done?

OpenStudy (anonymous):

\[f^{-1}(x) = \frac{ 5x-3 }{ 2 }\] now quite, now plug in 3 for x and that will give you\[f^{-1}(3)\]

jagr2713 (jagr2713):

5(3)-3/2 15-3/2 12/2 6

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

Good job!

jagr2713 (jagr2713):

so am done :D thxxxx

OpenStudy (anonymous):

Yes, now you're done

jagr2713 (jagr2713):

wait i got 6 for all of them

OpenStudy (anonymous):

What do you mean?

jagr2713 (jagr2713):

Which of the following statements have the same result? Explain each step in solving each one. f(1) when f(x) = 5x + 1 f-1(3) when f(x) = 2x plus 3, all over 5 3y - 7 = y + 5

jagr2713 (jagr2713):

i got 6 for all 3

OpenStudy (anonymous):

That's fine

OpenStudy (anonymous):

Explain your reasoning for all of them then :)

jagr2713 (jagr2713):

oh ok thx

jagr2713 (jagr2713):

wait you want to explain it now

OpenStudy (anonymous):

Just to make sure we did the right problem, your question was \[f(x) = \frac{ 2x+3 }{ 5 }\] and not \[f(x) = 2x+\frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

The explaining is your job :)

jagr2713 (jagr2713):

ok thx